OpenStudy (eujc21):

If the input current is of the form I_N sin(wt), of a parallel RLC circuit, determine the steady-state output voltage and the steady-state inductor current.

OpenStudy (eujc21):

So far just for the source being "IN" for voltage I have \[\frac{ dv_{out}^2 }{ dt^2 } + \frac{1}{RC}\frac{dV_{out}}{dt} +\frac{V_{out}}{R} - \frac{1}{C}\frac{dI_N}{dt} = 0\]

OpenStudy (eujc21):

When given \[I_Nsin(\omega t) = (I_C + I_L + I_R)\sin(\omega t)\] I get stuck because I am unsure if I am suppose to derive for \[[I_Nsin(\omega t)]\]

OpenStudy (eujc21):